OpenStudy (please_help):

How do you know if a quadratic equation will have one, two, or no solutions? How do you find a quadratic equation if you are only given the solution? Is it possible to have different quadratic equations with the same solution?

OpenStudy (anonymous):

1) put the equation as \[ax^2+bx+c=0\]then \[b^2-4ac\]say it all if \[b^2-4ac<0\] i.e. is negative, then there is no real solution if \[b^2-4ac=0\]then there is one solution and if \[b^2-4ac>0\] i.e. it is positive, then there are two solutions

OpenStudy (anonymous):

2) suppose you are given the solutions as \[r_1,r_2\] then you can find an equation that has these solutions by writing \[(x-r_1)(x-r_2)=0\] for example, if the two solutions are -3 and 5 you would write \[(x+3)(x-5)=0\] or \[x^2+2x-15=0\]

OpenStudy (anonymous):

typo there, should have been \[x^2-2x-15=0\] sorry

OpenStudy (anonymous):

yes, you can have two different quadratic equations that have the same solution, but one will be a multiple of the other. for example \[x^2-2x-15=0\] has solutions -3 and 5, but so does \[-2x^2+4x+30=0\] (i multiplied by -1) two equations with the same solution are called "equivalent"

OpenStudy (please_help):

Thank you so much!

OpenStudy (anonymous):

yw